# Efficient Neural Video Representation via Structure-Preseving Patch Decoding

### v1

@@ V Conclusion
We propose a neural video representation method based on structure-preserving patches (SPPs). In our approach, each video frame of a video is rearranged into spatially consistent patch images, enabling reconstruction that preserves global structure. This strategy mitigates the degradation commonly seen in conventional patch-based decoding. To facilitate smooth reconstruction, we designed a global-to-local decoder architecture. Early decoding layers are conditioned on the temporal index to model the global structure, while later layers incorporate patch-specific information to enhance local details. This separation allows the network to balance structural coherence and fine-grained fidelity effectively. Experimental results demonstrate that our method achieves higher reconstruction quality than existing NeRV-based baselines. These findings highlight the potential of structure-aware patch design for improving the fidelity and efficiency of neural video representations. In future work, we plan to explore more flexible adaptive patch layouts to further enhance spatial coherence and compression performance.

### v2

@@ V Conclusion
We propose a neural video representation method based on Structure-Preserving Patches (SPPs). In our approach, each video frame is split into spatially consistent patch images that maintain the original spatial structure. This strategy mitigates the degradation commonly seen in conventional methods and contribute to the fitting of global structures. To facilitate smooth reconstruction, we designed a global-to-local decoder architecture. Early decoding layers are conditioned on the temporal index to model the global structure, while later layers incorporate patch-specific information to enhance local details. This separation allows the network to balance structural coherence and fine-grained fidelity effectively. Experimental results demonstrate that our method achieves higher reconstruction quality than existing NeRV-based baselines. These findings highlight the potential of structure-aware patch design for improving the fidelity and efficiency of neural video representations. In future work, we plan to explore adaptive, content-aware patch layouts that dynamically adjust to scene complexity, aiming to further improve spatial coherence and compression performance.